Write the quadratic function in standard form and find the vertex.
f(x)=1/2x^2+10x+4

Answer by Boreal(15235)

Move the 4 to the right so (1/2)x^2+10x=-4
Factor out a (1/2)
(1/2)(x^2+20x)=-4
complete the square
(1/2)(x+10)^2=-4+50; the 10^2=100 is multiplied by (1/2)
(1/2)(x+10)^2-46=0
vertex is (-10, -46)
Check
vertex x is -b/2a=-10/1=-10
f(-10)=-46